OpenStudy (anonymous):

Any idea how to figure this out? :( Is there an equation or? Assume that it takes 10,000 Joules of energy to raise 100 grams of cake batter from room temperature to 90∘C (fully cooked). How many seconds would it take for a 100-Watt light bulb to output this much energy? (Hint: 1 Watt = 1 Joule per second)

OpenStudy (anonymous):

\[Power~[Watts]=\frac{ Energy~[Joules] }{ time~[seconds] }\] \[1~W = \frac{ 10000~J }{ t }\] Solve for t

OpenStudy (anonymous):

oops 100 W \[100~W=\frac{ 10000~J }{ t }\]

OpenStudy (anonymous):

Thank you!